OpenStudy (anonymous):

Identify the 8th term of the geometric sequence -2, -12, -72 ...

OpenStudy (anonymous):

a) -3,359,232 b) -559,872 c) 559,872 d) 3,359,232

OpenStudy (anonymous):

the rule is this ar^0+ar^1+ ar^7 1 2 8

OpenStudy (anonymous):

you your first term ar^0=-2 however r^0=1 so you ogt a*1=-2 or just a=-2 --- knowing a=-2 we plug that into your 2nd term to find r -2r=-12 r=6 *note that r^1=r --- we no know a and r and apply to the 8th term -2*6^7=-559872